OpenStudy (anonymous):

Pedro has created the function f(x) = the quantity of 4x minus 3, divided by 2 to represent the number of assignments he has completed, where x represents the number of weeks in the course. Pedro discovers that using the inverse function to solve for x = 30, he can predict when he will have 30 assignments completed. Explain to Pedro how to accomplish this, using complete sentences.

OpenStudy (anonymous):

\[4x - 3 \over 2\]

OpenStudy (rsadhvika):

Since \(\large f(x) = \frac{4x-3}{2}\) tells you # of assignments Pedro completes, given the week number, its inverse function should tell him how many weeks it should take to complete 30 assignments.

OpenStudy (rsadhvika):

Start by finding the inverse function of \(f(x)\)

OpenStudy (anonymous):

so were looking for f(-1)?

OpenStudy (rsadhvika):

yup, we're looking for \(f^{-1}(x)\)

OpenStudy (anonymous):

is this the equation where we input -1 for x?

OpenStudy (rsadhvika):

Nope, first we need to find the inverse funciton :- \(\large f(x) =y = \frac{4x-3}{2}\) \(\large y = \frac{4x-3}{2}\) multiply 2 both sides \(\large 2y = 4x-3\) add 3 both sides \(\large 2y+3 = 4x\) divide 4 both sides \(\large \frac{2y+3 }{4}= x\) switch x, y \(\large \frac{2x+3 }{4}= y\)

OpenStudy (rsadhvika):

thats the inverse function of f(x)

OpenStudy (anonymous):

Oh ok, thats when you switch x and y

OpenStudy (rsadhvika):

put x = 30 in this function, it will tell u how many weeks it takes to complete 30 assignments :- \(\large \frac{2x+3 }{4}= y\) put x = 30 \(\large \frac{2(30)+3 }{4}= y\) \(\large \frac{60+3 }{4}= y\) \(\large \frac{63 }{4}= y\) \(\large 15.75 = y\) so it wud take 16 weeks to complete 30 assignments !

OpenStudy (rsadhvika):

see if that makes some sense

OpenStudy (anonymous):

I think I got it

OpenStudy (anonymous):

Gimme a sec to write it down :)

OpenStudy (rsadhvika):

okie

OpenStudy (anonymous):

got it, thanks :)